### v1

# Debias Can be Unreliable: Mitigating Bias Issue in Evaluating Debiasing Recommendation

Abstract
Recent work has improved recommendation models remarkably by equipping them with debiasing methods. Due to the unavailability of fully-exposed datasets, most existing approaches resort to randomly-exposed datasets as a proxy for evaluating debiased models, employing traditional evaluation scheme to represent the recommendation performance. However, in this study, we reveal that traditional evaluation scheme is not suitable for randomly-exposed datasets, leading to inconsistency between the Recall performance obtained using randomly-exposed datasets and that obtained using fully-exposed datasets. Such inconsistency indicates the potential unreliability of experiment conclusions on previous debiasing techniques and calls for unbiased Recall evaluation using randomly-exposed datasets. To bridge the gap, we propose the Unbiased Recall Evaluation (URE) scheme, which adjusts the utilization of randomly-exposed datasets to unbiasedly estimate the true Recall performance on fully-exposed datasets. We provide theoretical evidence to demonstrate the rationality of URE and perform extensive experiments on real-world datasets to validate its soundness.
1 Introduction
Recommendation models suffer from various biases such as selection bias (Marlin et al. 2007; Hemminki and Sihvo 1993; Huang, Oosterhuis, and de Rijke 2022; Vernade and Cappé 2015) and popularity bias (Abdollahpouri and Mansoury 2020; Zhang et al. 2021; Zhu et al. 2021; Abdollahpouri et al. 2021), leading to deviations from true user preferences even harmful impacts to Web services (Chen et al. 2023; Klimashevskaia et al. 2023; Klimashevskaia 2022; Kowald, Schedl, and Lex 2020), such as amplifying the long-tail effect and hurting user trust (Abdollahpouri, Burke, and Mobasher 2017; Abdollahpouri et al. 2019; Abdollahpouri 2019). Undoubtedly, user’s true preferences can be accurately measured by the fully-exposed dataset which includes users’ feedback (e.g., ratings) across all items. The gold standard for evaluating debiasing techniques is thus utilizing such a fully-exposed dataset to assess the performance of recommendations (Gao et al. 2022a). However, due to the unavailability of fully-exposed datasets in practice (Gao et al. 2022b), randomly-exposed datasets, which contain users’ feedback on some randomly selected items, are widely employed for debiasing evaluation (Liu et al. 2020; Yuan et al. 2019; Rosenfeld, Mansour, and Yom-Tov 2017; Bonner and Vasile 2018; Schnabel et al. 2016; Wang et al. 2019; Ai et al. 2018; Chen et al. 2021). Notably, the amount of feedback in randomly-exposed datasets is also extremely limited, because randomly exposing items to collect user feedback is highly expensive, and the collection process can severely damage a platform’s revenue.
Figure 1: Preliminary analysis on KuaiRec, an only existing fully-exposed dataset, validating the inconsistency between Recall@$\overline{K}$ on the randomly-exposed data ($l$ = 80 samples per user) and Recall@$K$ on the fully-exposed data ($L$ = 3327 samples per user) when $K$ is small. We rank four AutoDebias models with different hyperparameters w.r.t. Recall@$\overline{5}$ and Recall@$K$, respectively. Notably, $\overline{K}\in$ [1,2,…$l$], $K\in$ [1,2,…$L$], where $l<<L$.
Previous work calculate widely-used Recall metric on randomly-exposed datasets in the same way as on fully-exposed datasets, which we refer to as the traditional evaluation scheme in subsequent descriptions. Nevertheless, we claim that traditional evaluation scheme is not suitable for randomly-exposed datasets when evaluating debiasing recommendation. Specifically, leveraging the only existing full-exposed dataset, KuaiRec (Gao et al. 2022b), our preliminary analysis reveals significant disparities between the evaluation results obtained from the randomly-exposed dataset and those derived from the fully-exposed dataset. Figure 1 is a quick example that presents the rankings of different models (i.e., AutoDebias models with different hyperparameters) based on the Recall@$\overline{5}$ on the randomly-exposed dataset and Recall@$K$ on the fully-exposed dataset. As can be seen, the model ranking deduced from Recall@$\overline{5}$ is different from those derived from Recall@$K$ when $K\leq 100$. Given that user typically browse only a limited number of items during an online deployment stage, Recall@$K$ with a small value of $K$ (e.g., $K=50$) on the fully-exposed dataset represents desired debiasing quality and correct model ranking. Consequently, the conclusion derived from the traditional evaluation scheme on the randomly-exposed dataset such as Recall@$\overline{5}$ might be unconvincing.
To comprehensively investigate the underlying factors contributing to the inconsistency between the Recall@$\overline{K}$ on the randomly-exposed dataset and Recall@$K$ on the fully-exposed dataset, we undertake a theoretical analysis (cf. Theorem 3) of the evaluation performed on a randomly-exposed dataset. Our analysis demonstrates that the Recall@$\overline{K}$ on the randomly-exposed dataset (e.g., Recall@$\overline{5}$) has a weak correlation with the Recall@$K$ using a small value of $K$ on the fully-exposed dataset (e.g., Recall@50). Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al. 2023). In addition, we conduct extensive experiments on real-world datasets to further validate the variance of such correlation (cf. Figure 1).
Given that the traditional evaluation scheme on the randomly-exposed dataset has limited capacity to reflect the desired Recall@$K$ performance on the fully-exposed dataset, it is inadequate to conclude whether previous debiasing methods have effectively mitigated biases. The observed performance gain of debiasing methods on the traditional evaluation scheme can be attributed to optimizing Recall@$K$ under a large value of $K$ on the fully-exposed dataset using techniques that lack genuine debiasing effects. For instance, hard negative sampling that theoretically optimizes Recall@$K$ with a large $K$ on fully-exposed datasets (Shi et al. 2023), can outperform the representative debiasing method (Chen et al. 2021) under the traditional evaluation scheme (see Table 3). Considering the cost of constructing fully-exposed dataset, unbiasedly estimating the Recall@$K$ performance with a small value of $K$ on the fully-exposed dataset from randomly-exposed datasets is crucial for the evaluation of debiasing methods.
Figure 2: Illustration of URE. In randomly-exposed data, $n$ is the number of positive items with existing feedback and $m$ is the number of positive items with existing feedback ranked within top-$K$ of all items.
Towards this goal, we propose the Unbiased Recall Evaluation (URE) scheme as illustrated in Figure 2. We set the target as estimating the Recall@$K$ on fully-exposed data, i.e., the ratio of items with positive feedback ranked before position $K$, denoted as $\frac{M}{N}$. Considering that a randomly-exposed dataset only includes user feedback on randomly selected items, we can treat it as a representative subset of the fully-exposed dataset, wherein both positive and negative feedback are missing at random. In URE, we sort the model’s prediction scores of all candidate items and utilize the score of $(K+1)$-th item as the threshold to calculate the positive ratio on the randomly-exposed dataset, where the positive ratio is the proportion of items with positive feedback in the randomly-exposed dataset that rank before the threshold. Hence such a positive ratio can be used to unbiasedly estimate $\frac{M}{N}$ on fully-exposed dataset. This is due to the models’ prediction information of all candidate items, which is not considered by traditional evaluation scheme. Through theoretical analysis and extensive experiments, we validate that URE can unbiasedly estimate the Recall@$K$ on fully-exposed data with a desired value of $K$.
In this work, our contributions are threefold.
  • We reveal the inconsistency between the widely adopted traditional evaluation scheme on the randomly-exposed dataset and fully-exposed dataset (the gold standard evaluation), questioning empirical conclusions on existing debiasing methods in previous studies.
  • We theoretically and empirically analyze the relation between the Recall@$\overline{K}$ on the randomly-exposed dataset and the Recall@$K$ on the fully-exposed dataset.
  • We propose the URE scheme to unbiasedly estimate the target Recall@$K$ performance with randomly-exposed data, and conduct extensive theoretical and empirical analysis, validating the effectiveness of the URE scheme.
2 Preliminary
Table 1: Notations and definitions.
Notations Annotations
$D_{full}$ A fully-exposed dataset
$D_{rand}$ A randomly-exposed dataset
$L/l$ The amount of feedback from a single user
$K$ Top-K on $D_{full}$, $K\in$ [0,1,…$L$]
$\overline{K}$ Top-K on $D_{rand}$, $\overline{K}\in$ [0,1,…$l$] ($l<<L$)
Recall@$K$ Recall metric at position $K$ on $D_{full}$
Recall@$\overline{K}$ Recall metric at position $\overline{K}$ on $D_{rand}$
$\widehat{\text{Recall@}K}$ The estimation of $\text{Recall}@K$ by URE scheme
Given a user ($u$) set and an item ($i$) set, fully-exposed dataset $D_{full}$ includes the user feedback $r_{u,i}$ (e.g., ratings and clicks) on all candidate items, which is usually unavailable in practice. The only existing $D_{full}$ for research is KuaiRec (Gao et al. 2022b). All candidate items can be categorized into items with positive feedback that user likes ($r_{u,i}=1$) and items with negative feedback ($r_{u,i}=0$) based on user’s true interests. For the evaluation of debiasing recommendation, researchers mostly resort to randomly-exposed datasets $D_{rand}$ as a proxy to evaluate debiasing methods. However, randomly-exposed dataset $D_{rand}$ collects the user feedback on some randomly selected items ($\in$ all candidate items) through a costly randomized intervention experiment (Liu et al. 2020; Yuan et al. 2019; Rosenfeld, Mansour, and Yom-Tov 2017; Bonner and Vasile 2018; Schnabel et al. 2016; Wang et al. 2019; Ai et al. 2018; Chen et al. 2021), i.e., a large amount of user feedback is ramdomly missing. During the evaluation stage, recommendation models ($\hat{y}=M(u,i)$) rank the items with feedback based on $\hat{y}$ to generate a personalized ranking list $Rank_{u}$ for each user. In traditional evaluation scheme on $D_{full}$, we can obtain the true performance of debiased recommendation models by calculating the Recall@$K$. Notably, The Recall metric is the most important metric in the candidate generation stage of industrial recommendation systems. It is also a commonly used metric in research on debiasing recommendations. Recall@$K$ on $D_{full}$ is obtained by the equation 1.
$\text{Recall}@K=\frac{|{k\in Rank_{u}+:k\leq K}|}{|Rank_{u}+|},K\in[1,2,...,L],$ (1)
where the $Rank_{u}+$ collects the positions of all items with positive feedback within $Rank_{u}$.
Unfortunately, as mentioned above, fully-exposed dataset is unavailable on most recommendation platforms. The Recall@$K$ in the equation 1 is not supported if the feedback of any candidate item is missing. For the evaluation of debiasing recommendation, researchers employ traditional evaluation scheme on $D_{rand}$ to calculate the Recall performance, denoted as Recall@$\overline{K}$.
Figure 3: The Pearson correlation coefficients or Kendall coefficients between Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$. We highlight the value of $K$ when each curve reaches its maximum correlation coefficient as $K_{max}$. (a) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5. (b) The effect of $\overline{K}$, where $\overline{N}$ is fixed to 80. (c) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5. The correlations are statistically significant with $p$ $<$ 0.01.
$\begin{split}\text{Recall}@\overline{K}=\frac{|{k\in\overline{Rank_{u}}+:k\leq\overline{K}}|}{|\overline{Rank_{u}}+|},\\ \overline{K}\in[1,2,...,l],\quad l<<L,\end{split}$ (2)
where $\overline{Rank_{u}}+$ contains the positions of items with positive feedback of user $u$ in $\overline{Rank_{u}}$. However, in this work, we find that Recall@$\overline{K}$ on the randomly-exposed dataset is inconsistent with Recall@$K$ on fully-exposed dataset. Therefore traditional evaluation scheme is not suitable for randomly-exposed datasets. We will detail the inconsistency and unreliability of such a traditional evaluation scheme in the following sections.
[…]
In this section, we first conduct theoretical analysis to discover why the traditional evaluation scheme on randomly-exposed data is unreliable by inspecting the relationship between Recall@$K$ metric on fully-exposed data and Recall@$\overline{K}$ metric on randomly-exposed data. Then, we substantiate their correlation through empirical validation.
3.1 Theoretical Guarantee
In this subsection, we give the theoretical derivation of the relationship between the Recall@$K$ on $D_{full}$ and the Recall@$\overline{K}$ on $D_{rand}$.
Theorem 1.
Assuming that we have $N^{+}$ positive samples and $N^{-}$ negative samples on $D_{full}$, with a total sample size of $N=N^{+}+N^{-}$, our $D_{rand}$ samples $\overline{N}$ samples from $D_{full}$. We denote the set of all combinations of $D_{rand}$ of size $\overline{N}$ as $\mathcal{S}_{\overline{N}}$ and the set of all permutaions of $D_{full}$ as $\mathcal{R}$. Then, when $K$ on $D_{full}$ and $\overline{K}$ on $D_{rand}$ satisfy $\overline{K}=\frac{\overline{N}}{N}\cdot K$, we have:
$E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]-\text{Recall@}K\right]=0,$ (3)
where $E$ denotes the Expectation funtion.
The proof of Theorem 3 can be found in Appendix A.2. From Theorem 3, we can observe that the Recall@$\overline{K}$ on $D_{rand}$ will exhibit a strong correlation with the Recall@$K$ on $D_{full}$ if the value of $K$ is around $\frac{{N}}{\overline{N}}\cdot\overline{K}$. It should be noted that $\overline{N}$ is much smaller than $N$ since the randomly-exposed dataset is a sparsely sampled subset of the fully-exposed dataset. Therefore, given a specific value of $\overline{K}$, the Recall@$\overline{K}$ on $D_{rand}$ only exhibits a strong correlation with the Recall@$K$ with much larger $K$ values on $D_{full}$. However, given that user typically browse only a limited number of items during an online deployment stage, Recall@$K$ with a small value of $K$ (e.g., $K=50$) on the $D_{full}$ represents desired debiasing quality and correct model ranking. Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al. 2023).
3.2 Empirical Results
In this subsection, we conduct empirical experiments to validate the relationship between Recall@$K$ metric on fully-exposed data and Recall@$\overline{K}$ metric on randomly-exposed data.
We train multiple AutoDebias (Chen et al. 2021) models w.r.t. different hyperparameters and evaluate their Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$. Then we calculate Pearson and Kendall coefficients between Recall@$\overline{K}$ and Recall@$K$ to measure their linear correlation and ranking correlation (Abdi 2007), respectively. The results are shown in Figure 3. We highlight the value of $K$ when each curve reaches its maximum correlation coefficient as $K_{max}$. From the Figure 3, we have the following observations:
  • When $\overline{N}$ and $\overline{K}$ are hold constant, the highest correlation coefficients between Recall@$\overline{K}$ and Recall@$K$ are achieved with large values of $K$. As depicted in Figure 3a and Figure 3b, the Pearson correlation coefficient between Recall@$\overline{K}$ and Recall@$K$ surpasses 0.9 for values of $K$ exceeding 100, while it remains below 0.6 for values of $K$ below 10. This is consistent with our observation from Theorem 3 that the Recall@$\overline{K}$ on $D_{rand}$ only exhibits a strong correlation with the Recall@$K$ with large values of $K$ on $D_{full}$. Hence, the evaluation of models’ Recall@$\overline{K}$ performance on $D_{rand}$ may not adequately reflect Recall@$K$ performance with small $K$ on $D_{full}$.
  • When fixing $\overline{K}$, the value of $K_{max}$ decreases as the sample number $\overline{N}$ increases. As shown in Figure 3a and Figure 3c, the inverse relationship between $K_{max}$ and $\overline{N}$ is consistent with the conclusion in Eq. (3) that $K=\frac{N}{\overline{N}}\cdot\overline{K}$. This is intuitively understandable since an increase in the sample size $\overline{N}$ of $D_{rand}$ leads to a stronger and more reliable representation of the fully-exposed dataset $D_{full}$. However, in practice, scaling up the sample size is often impractical due to substantial economic constraints.
  • When fixing $\overline{N}$, the value of $K_{max}$ decreases as the $\overline{K}$ decreases. As shown in Figure 3b, the positive correlation between $K_{max}$ and $\overline{K}$ is also consistent with the conclusion in Eq. (3) that $\overline{K}=\frac{\overline{N}}{N}\cdot K$. However, as we discussed before, due to the frequent occurrence of $\overline{N}\ll N$, even when considering Recall@$\overline{K}$ with $\overline{K}=1$, it is still not possible to unbiasedly and accurately estimate Recall@$K$ for small values of $K$.
  • When fixing $\overline{K}$, the highest Kendall correlation coefficient increase as sample number $\overline{N}$ increase. Figure 3c shows that when sample number $\overline{N}$ is small, Kendall correlation coefficient will be small for any Recall@$K$. Low Kendall correlation coefficient means that if model A has a higher Recall@$\overline{K}$ value on $D_{rand}$ than model B, it does not necessarily indicate that A performs better than B with Recall@$K$ on $D_{full}$. Therefore, it is unreliable to use the Recall@$\overline{K}$ on $D_{rand}$ for model selection.
In conclusion, we prove from both theoretical and experimental aspects that there is a relationship between Recall@$K$ on the fully-exposed dataset and Recall@$\overline{K}$ on the randomly-exposed dataset. Specifically, Recall@$\overline{K}$ is associated with Recall@$K$ when the value of $K$ is relatively large. Given that the gold standard for debiasing recommendation desires Recall@$K$ with small values of $K$ (e.g., $K=50$), using traditional evaluation scheme on randomly-exposed dataset directly is problematic.
4 Entrapment Study
[…]
4.1 Hard Negative Sampling
In this subsection, we begin by introducing the concept of hard negative sampling. Subsequently, we present a theorem delineating the relationship between hard negative sampling and the optimization of Recall@$K$ on $D_{full}$ with the desired value of $K$.
Hard negative sampling refers to oversampling high-scored negative items for training. Taking the representative method DNS($H$, $U$) (Shi et al. 2023) as an example, it first uniformly selects $U$ negative samples and then selects the top $H$ hard samples with the highest prediction scores by the model for training. Particularly, as $H/U$ decreases, it implies that the selected hard negative samples become more challenging for the model. We thus define $H/U$ as the hardness of hard negative sampling.
We can present a theorem elucidating the relationship between hard negative sampling and optimizing Recall@$K$ on $D_{full}$.
Theorem 2.
When the error function $\delta$ is the BCE loss function, the model equipped with hard negative sampling DNS($H$, $U$) approximately optimizes Recall@$K$ metric. Meanwhile, the larger $H/U$ is, the Recall@$K$ metric with a larger $K$ on $D_{full}$ is optimized.
The proof can be found in Appendix A.3. Based on the Theorem 2, we conclude that the hard negative sampling approach could optimize Recall@$K$ with large values of $K$ on $D_{full}$ by using large $H/U$.
Table 2: The statistics of datasets.
Dataset ##\##Users ##\##Items Biased train test
Yahoo!R3 15,400 1,000 311,704 48,600 ($D_{rand}$)
KuaiRec 1411 3327 1,934,404 2,742,166 ($D_{full}$)
Table 3: The comparison between NegS and representative methods with debiasing techniques under the traditional evaluation scheme on $D_{rand}$. The best results are in bold and the second best are underlined. The results are statistically significant with $p<$ 0.01.
recall@$\overline{5}$ MF LightGCN
Yahoo!R3 KuaiRec Yahoo!R3 KuaiRec
Bias 0.7085 0.3546 0.7883 0.4859
IPS 0.6572 0.3102 0.7584 0.4889
DR 0.7195 0.5465 0.7576 0.4849
AutoDebias 0.8085 0.5482 0.8046 0.5181
NegS 0.8330 0.5358 0.8141 0.5288
[…]
4.2.1 Experimental Setup.
To assess the reliability of the traditional evaluation scheme on $D_{rand}$, we use two widely-used datasets for our following experiments: Yahoo!R3 and KuaiRec. The statistics related to these datasets can be found in Table 2. The KuaiRec is the only existing full-exposed dataset that also can be used to construct randomly-exposed set and training set. However, Yahoo!R3 only consists of a biased set for training and an randomly-exposed set for testing. We utilize two representative recommendation models, Matrix Factorization (MF) (Koren, Bell, and Volinsky 2009) and graph neural network model LightGCN (He et al. 2020), as the backbones. We compare NegS with following debiasing methods: IPS (Schnabel et al. 2016), DR (Wang et al. 2019), AutoDebias (Chen et al. 2021). All the models are optimized by Adam optimizer. AutoDebias uses random search to find the best hyperparameters while other models use grid search. The search space of learning rate and weight decay are [1e-4, 1e-3, 1e-2, 1e-1]. For the proposed NegS, we tune $H/U$ and $U$ within the ranges of [0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1] and [50,100,200] respectively. The detialed setup can be found in Appendix A.7.
4.2.2 Observations.
We compare the NegS to debiasing baselines using the traditional evaluation scheme on $D_{rand}$ in Table 3. For NegS, we utilize a large $H/U$ to obtain the best results. As shown in Table 3, we highlight that the NegS exhibits comparable or even better performance compared to the debiasing models on Recall@$\overline{K}$ metric on two datasets. From this example, we can infer that a model that optimizes the Recall@$K$ metric with a large $K$ on $D_{full}$, even without any debiasing technologies, can also achieve remarkable performance under the traditional evaluation scheme on $D_{rand}$. This shows that the traditional evaluation scheme on $D_{rand}$ cannot accurately identify effective debiasing models. As a result, the conclusions drawn from evaluating debiasing techniques in this manner may be unreliable, necessitating significant effort in future work for further investigation.
5 URE scheme
Figure 4: Illustration of the proposed URE scheme on unbiasedly estimating Recall@$K$ on $D_{full}$.
Considering the impractical accessibility of $D_{full}$, the pivotal role of $D_{rand}$ emerges in enabling an unbiased estimation of Recall@$K$ on $D_{full}$. Nevertheless, as mentioned above, our contention is that the traditional evaluation scheme on $D_{rand}$ exhibits bias. In response to this challenge, we introduce a pioneering evaluation scheme coined URE, delineated in Figure 4. URE relies solely on $D_{rand}$ to achieve an unbiased estimation of Recall@$K$ on $D_{full}$, which consists of the following steps:
  • Obtain the predictions $\hat{y}$ ($\hat{y}=M(uid,iid)$) for all candidate items ($iid$$\in$[1, …, N]), where the $M$ is the model to be tested.
  • Sort these candidate items ($iid$$\in$[1, …, N]) in descending order by their predictions $\hat{y}$, denoted as Sorted ItemID in Figure 4.
  • Assign the labels (i.e. feedback) in $D_{rand}$ (positive : 1, negative : 0, Missing labels : \) to each candidate item in Sorted ItemID, denoted as Sorted $D_{rand}$ in Figure 4.
  • Locate the ($K$+1)-th item in Sorted ItemID and its corresponding prediction $\hat{y}$.
  • Compute the $\widehat{\text{Recall@}K}$ for this user:
    $\widehat{\text{Recall@}K}=\frac{m}{n},$
    where $n$ denotes the number of items with positive label in Sorted $D_{rand}$, and $m$ denotes the number of these items with higher $\hat{y}$ than the ($K$+1)-th item.
  • We report the averaged $\widehat{\text{Recall@}K}$ across all users as an unbiased estimate of the model’s overall Recall@$K$ performance on $D_{full}$.
Theoretical Guarantee. The following theorem demonstrates that the estimates provided by URE are unbiased.
Theorem 3.
(URE Unbiasness) For all given $K$, we have:
$E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall@}K}\right]=\text{Recall}@K,$ (4)
where $\mathcal{S}_{\overline{N}}$ denotes the set of all possible random samplings of size $\overline{N}$.
The proof can be found in Appendix A.4. This signifies that the estimates derived from URE are unbiased.
In addition to the theoretical analysis, we provide an intuitive explanation to show how our URE can unbiasedly estimate the Recall@$K$ on $D_{full}$. The core reliance is that $D_{rand}$ represents a random selection from the sorted list of $D_{full}$. Hence, the ratio of items with positive label in $D_{rand}$ which have a higher $\hat{y}$ than the ($K$+1)-th item can be used to unbiasedly estimate the positive ratio of $D_{full}$. Our proposed URE scheme, despite its apparent simplicity, can generate outstanding predictive accuracy. This is due to the models’ prediction information of all candidate items, which is not considered by traditional evaluation scheme. We will demonstrate the effectiveness of this scheme in following sections.
Figure 5: (a) The Kendall rank coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$. (b) The Kendall rank coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$. We mark $K_{max}$ when every curve reaches its highest coefficient. The rank correlations in Figure (a) and (b) are statistically significant with $p<$ 0.01.
6 Experiments
Table 4: The evaluation results by using the URE scheme. We also show the Recall@$\overline{5}$ under the traditional evaluation scheme for comparison. The best results are in bold and the second best are underlined. Notably, the Recall@$5$ metric is not available in Yahoo!R3, as it lacks a fully-exposed dataset. The results are statistically significant with $p<$ 0.01.
Backbone Model KuaiRec Yahoo!R3
Traditional scheme Fully-Exposed Dataset URE scheme Fully-Exposed Dataset Traditional scheme URE scheme
Recall@$\overline{5}$ Recall@$5$ $\widehat{\text{Recall@5}}$ $\widehat{\text{Recall@500}}$ Recall@500 Recall@$\overline{5}$ $\widehat{\text{Recall@5}}$ $\widehat{\text{Recall@500}}$
MF Bias 0.3546 0.0215 0.0213 0.3280 0.3237 0.7085 0.0131 0.6777
IPS 0.3102 0.0213 0.0230 0.2771 0.2743 0.6572 0.0095 0.6236
DR 0.5465 0.0353 0.0362 0.5077 0.5050 0.7195 0.0075 0.6879
AutoDebias 0.5482 0.0355 0.0352 0.5129 0.5097 0.8085 0.0550 0.7849
NegS 0.5358 0.0327 0.0326 0.5052 0.5012 0.8330 0.0490 0.8001
LightGCN Bias 0.4859 0.0257 0.0267 0.4503 0.4482 0.7883 0.0491 0.7680
IPS 0.4889 0.0266 0.0273 0.4539 0.4522 0.7584 0.0457 0.7367
DR 0.4849 0.0262 0.0267 0.4500 0.4457 0.7576 0.0455 0.7373
AutoDebias 0.5181 0.0317 0.0323 0.4571 0.4558 0.8046 0.0592 0.7820
NegS 0.5288 0.0237 0.0245 0.4954 0.4944 0.8141 0.0616 0.7913
In this section, we conduct experiments to verify the effectiveness of the URE scheme and re-evaluate the debiasing methods using the proposed URE scheme.
6.1 Unbiasedness Analysis
In this subsection, we test the AutoDebias model using the KuaiRec dataset to verify that the proposed $\widehat{\text{Recall@}K}$ in the URE scheme is an unbiased estimation of Recall@$K$ on $D_{full}$.
In Figure 5a, we investigate whether there is order preservation between Recall@$K$ and our $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$. Concretely, we randomly select model’s hyperparameter combinations to train the AutoDebias model on the Kuairec dataset. Then we calculate the Kendall rank correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$, as well as Recall@$\overline{K}$ to demonstrate the ranking consistency. As shown in Figure 5a, when $K$ is small (e.g. $K=30$), both Recall@$\overline{1}$ and Recall@$\overline{5}$ exhibits poor correlations with Recall@$K$. In contrast, the Kendall correlation coefficient between the estimated $\widehat{\text{Recall@}K}$ and Recall@$K$ is higher than 0.9. This means that the model we choose based on $\widehat{\text{Recall@}K}$ is more likely to perform better on Recall@$K$ on $D_{full}$, which further demonstrates the effectiveness of the URE scheme.
In Figure 5b, we report the Kendall rank correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$. As we desired, only when K is the same, Recall@$K$ and $\widehat{\text{Recall@}K}$ have the highest correlation coefficients. That is, URE allows us to select the best debiased model based on the Recall@$K$ with a desired $K$.
Based on our discussion, we show that our URE scheme unbiasedly estimates Recall@$K$ for different values of $K$. Moreover, the high Kendall coefficients between $\widehat{\text{Recall@}K}$ and Recall@$K$ ensure an accurate debiased model selection.
6.2 Re-evaluation Results
In this subsection, we re-evaluate some classical debiasing methods, and report the results in Table 4. We have the following key findings:
  • Based on the results on KuaiRec dataset, we observe that the evaluation result $\widehat{\text{Recall@5}}$ of the URE scheme aligns closely with the true Recall@5 on the fully-exposed dataset. This observation suggests a strong unbiasedness in the URE scheme. In contrast, the evaluation results of the traditional scheme exhibit a significant discrepancy from the true Recall@5, and it fails to ensure the preservation of order in the results.
  • In some instances, the $\widehat{\text{Recall@5}}$ estimated by the URE scheme is inversely ordered compared to the actual Recall@5. This phenomenon can occur when the actual Recall@5 values are closely aligned, as is the case with the MF-based DR method at 0.353 and the MF-based AutoDebias method at 0.355. We contend that such a degree of deviation is within an acceptable range.
  • When using $\widehat{\text{Recall@5}}$ to verify the debiasing performance, some debiasing methods perform worse than the Bias model, which encourages the development of further debiasing techniques.
  • Based on the proposed URE, the performance of NegS is sub-optimal as expected in most instances, which implies that URE possesses potential merits in selecting truly effective debiasing approaches. In the LightGCN-based method on Yahoo!R3, the Recall@5 of NegS exceeds that of all other methods, which we attribute to the fact that these methods are not successfully debiasing in this case.
  • There is an inconsistency in the ranking results based on $\widehat{\text{Recall@}K}$ metric with large and small $K$ values. This means we should develop different debiasing methods to accommodate Recall@$K$ with different $K$ on $D_{full}$.
7 Related Work
7.1 Debiasing Recommendation
Recent recommendation debiasing strategies can be divided in two main categories: data-level debiasing and representation-level debiasing. At the data level, debiasing is always achieved by inverse propensity reweighting  (Schnabel et al. 2016; Wang et al. 2016; Li et al. 2023c) and doubly robust learning (Dai et al. 2022; Li et al. 2023b; Li, Zheng, and Wu 2023). At the representation level, information bottleneck (Wang et al. 2020), causal representation learning (Liu et al. 2020; Yang et al. 2023), and invariant representation learning (Wang et al. 2022; Yang et al. 2023; Pan et al. 2023) are widely used. For evaluating the debiasing effectiveness of these methods, researchers often evaluate Recall@$\overline{K}$ performance (e.g., Recall@$\overline{5}$) of the learned recommendation models on publicly available randomly-exposed data (Chen et al. 2023) — Coat, Yahoo!R3 and KuaiRand. However, we point out such an evaluation scheme on $D_{rand}$ may be problematic.
7.2 Sampling Metrics
The traditional evaluation scheme that calculates Recall@$\overline{K}$ metric on $D_{rand}$ can be thought of as a type of sampling metric. (Krichene and Rendle 2020) is the first work to highlight the inconsistency between sampled metrics on the sampled data and exact metrics on full data. It also proposes a method based on the sampled data to unbiasedly estimated Recall@$K$ metrics on full data. The simple estimation method suffers from the problem of large variance. Subsequent work (Li et al. 2023a, 2020; Jin et al. 2021; Dallmann, Zoller, and Hotho 2021) improves the estimation scheme and reduce the estimation variance. However, all the works mentioned above only focus on the case where the test set contains only one positive sample. In this way, the conclusions are not applicable to the debiasing task in that the test set contains multiple positive samples. Furthermore, there are no discussions about the effect of traditional evaluation scheme on the debiasing task.
8 Conclusion And Future Work
In this paper, we reviewed traditional evaluation scheme on the randomly-exposed datasets by comparing it with the gold standard evaluation on the fully-exposed datasets. According to the theoretical and empirical analyses, we revealed the inconsistency between them w.r.t. the values of $\overline{K}$ and $K$. We conducted an entrapment study, where a hard negative sampling trick without any debiasing techniques can achieve incredible debiasing performance by using the traditional evaluation scheme on the randomly-exposed datasets, further validating the traditional evaluation scheme on the randomly-exposed datasets may not reflect true debiasing performance and results in unreliable conclusions. Accordingly, we designed the URE scheme to unbiasedly estimate the true Recall performance of the fully-exposed datasets. Both theoretical and empirical analyses have validated the effectiveness of our proposal.
In future work, we will investigate the unbiased estimation of other Top-$K$ metrics, like Precision and NDCG on the fully-exposed dataset to further extend our conclusions. Besides, improving URE from the perspective of bias-variance balancing is another interesting direction. We hope that the insight of this paper could bring new inspiration for debiasing recommendations.
Appendix A Appendix
A.1 Datasets
We use two real-world datasets (Yahoo!R3 and KuaiRec) for our experiments.
  • Yahoo!R3: This dataset is widely-used in debiasing literature which consists of a biased dataset and a randomly-exposed dataset collected by a random logging strategy. We follow (Chen et al. 2021) and randomly sample 10% of the randomly-exposed data for debiasing training (required by AutoDebias and IPS), and take the remaining 90% as $D_{rand}$ for evaluating. The interactions between users and items are ratings ranging from one to five. The ratings larger than three are labeled as positive ($r_{u,i}$ = 1), otherwise negative ($r_{u,i}$ = 0). Notably, the Recall@$K$ metric is not available in Yahoo!R3, as it lacks a fully-exposed dataset. However, in our proposed URE scheme, we can use $D_{rand}$ to compute $\widehat{\text{Recall@}K}$ for unbiased estimation of Recall@$K$.
  • KuaiRec: This dataset is the only existing fully-exposed dataset. We construct three types of datasets from this original fully-exposed data: biased data for training, fully-exposed data $D_{full}$ for evaluating, and randomly-exposed data $D_{rand}$ for evaluating. Concretely, we sample items for each user inversely proportional to their video duration and combine them with their ratings to obtain biased data. Interactions on the $k$-th shortest video are sampled with probability $p=e^{k/\lambda}/\sum_{k}e^{k/\lambda}$, $\lambda$ controls the degree of bias towards short videos. The remaining items and ratings is considered as the fully-exposed data ($D_{full}$) for evaluating. Following (Marlin and Zemel 2009), we construct randomly-exposed data ($D_{rand}$) by uniformly sampling $\overline{N}$ items for each user on $D_{full}$. The watch_ratio is binarized with threshold 2 following (Gao et al. 2022a). That is, the video watching ratio larger than 2 is labeled as positive ($r_{u,i}$ = 1), otherwise negative ($r_{u,i}$ = 0).
A.2 Proof of Theorem 3
Proof.
We first calculate the $E_{\mathcal{R}}[\text{Recall@}K]$. Since we consider all possible ranking, the possibility of the x positive samples contained in the top K is $\frac{C_{N^{+}}^{x}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$ where $C^{n}_{N}:=\frac{N!}{(N-n)!n!}$ is the number of combinations obtained by randomly sampling $n$ samples from $N$. Thus, we have:
$E_{\mathcal{R}}\left[\text{Recall}@K\right]=\sum_{x=1}^{min(K,N^{+})}\frac{x}{N^{+}}\frac{C^{x}_{N^{+}}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$ (5)
$=\sum_{x=1}^{min(K,N^{+})}\frac{xN^{+}K}{xN^{+}N}\frac{C^{x-1}_{N^{+}-1}C^{(K-1)-(x-1)}_{(N-1)-(N^{+}-1)}}{C_{N-1}^{K-1}}$ (6)
$=\frac{K}{N}\left[\sum_{x=1}^{min(K,N^{+})}\left[C_{N^{+}-1}^{x-1}C_{(N-1)-(N^{+}-1)}^{(K-1)-(x-1)}/C^{K-1}_{N-1}\right]\right]$ (7)
$=\frac{K}{N}\left[\ C_{N-1}^{K-1}/C_{N-1}^{K-1}\right]=\frac{K}{N}.$ (8)
As for the $E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]$, we have:
$E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]=E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right].$ (9)
Since we consider all possible permutations, $E_{\mathcal{R},\mathcal{S}_{\overline{N}}}\left[\text{Recall@}\overline{K}|{N^{+}}^{\prime}\right]$ can be transformed into the scenario of randomly selecting $\overline{K}$ samples from $\overline{N}$ samples ($N^{+}$ positive samples among these $N$ samples) and determining how many of the selected $\overline{K}$ samples are positive. The problem is similar to Eq 5. Thus we have:
$E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right]$ (10)
$=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x}{{N^{+}}^{\prime}}\frac{C^{x}_{{N^{+}}^{\prime}}C^{\overline{K}-x}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}}^{\overline{K}}}\right]$ (11)
$=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x\overline{K}{N^{+}}^{\prime}}{x{N^{+}}^{\prime}\overline{N}}\frac{C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{N-{N^{+}}^{\prime}}}{C_{N-1}^{\overline{K}-1}}\right]$ (12)
$=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\frac{\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}-1}^{\overline{K}-1}}\right]$ (13)
$=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\right]=\frac{\overline{K}}{\overline{N}}=\frac{K}{N}=E_{\mathcal{R}}[\text{Recall@}K].$ (14)
A.3 Proof of Theorem 2
Proof.
It has been proven that when optimizing the BPR loss function using uniform negative sampling, the optimization objective of the model is AUC in the previous study (Rendle et al. 2009). Therefore, we only need to prove that the BCE loss function commonly used in debiasing tasks is an upper bound of the BPR loss function. Specifically, the BCE loss function is:
$\delta_{BCE}=\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\left[-\log(\sigma(f_{\theta}(u,i))-\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}\log(1-\sigma(f_{\theta}(u,j))\right],$ (15)
where $\sigma(x)=1/(1+e^{-x})$, $f_{\theta}(u,i)$ denotes a model with trainable parameter $\theta$ for predicting the rating of item $i$ by user $u$, and $\mathcal{D}$ refers to the train dataset containing ($u$, $i$) pairs.
BPR loss function:
$\delta_{BPR}=\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log(\sigma(f_{\theta}(u,i)-f_{\theta}(u,j)),$ (16)
where $\mathcal{I}$ denotes the item set.
By transforming the BCE loss function, we can obtain:
$\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\left[-\log(\sigma(f_{\theta}(u,i))-\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}\log(1-\sigma(f_{\theta}(u,j))\right]$ (17)
$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log(\sigma(f_{\theta}(u,i)\cdot(1-\sigma(f_{\theta}(u,j)))))$ (18)
$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\left[\frac{1}{1+e^{-f_{\theta}(u,i)}}\cdot\frac{e^{-f_{\theta}(u,j)}}{1+e^{-f_{\theta}(u,j)}}\right]$ (19)
$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\frac{1}{1+e^{-f_{\theta}(u,i)}+e^{f_{\theta}(u,j)}+e^{-f_{\theta}(u,i)+f_{\theta}(u,j)}}$ (20)
$\geq$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\frac{1}{1+e^{-(f_{\theta}(u,i)-f_{\theta}(u,j))}}=\delta_{BPR}.$ (21)
Thus, we have proved that the commonly used BCE loss function is an upper bound of the BPR loss function. Therefore, optimizing the BCE loss function is equivalent to optimizing the upper bound of the BPR loss function, which equals optimizing AUC.
Based on the aforementioned mentioned equivalence and referring to Theorem 1 and Theorem 3 from (Shi et al. 2023), we derive that optimizing BCE loss enhanced with DNS($H$, $U$) approximates the optimization of the Recall@$K$ metric. Specifically, the larger the $H/U$ is, the Recall@$K$ metric with a larger $K$ on $D_{full}$ is optimized. ∎
A.4 Proof of Theorem 3
Proof.
Since Recall is only related to positive samples, we can only consider positive samples. Let’s assume that in a fully exposed scenario, there are a total of $N$ positive samples, out of which $M$ samples are ranked higher than the sample ranked at $K+1$ in Sorted ItemID in Figure 4. Assuming that we have randomly sampled $\overline{N}$ samples, which have $n$ positive samples and the rest are negatives. Since the sampling process is random, the probability of the number $m$ of positive samples that rank higher than the sample ranked at $K+1$ in Sorted ItemID can be calculated as $(C^{m}_{M}C^{n-m}_{N-M}/C^{n}_{N})$, thus we have:
$E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall}@K}\right]$ $=E_{n^{\prime}}\left[E_{\mathcal{S}_{\overline{N}}}[m/n|n^{\prime}]\right]$ (22)
$=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{m}{n^{\prime}}\frac{C^{m}_{M}C^{n^{\prime}-m}_{N-M}}{C^{n^{\prime}}_{N}}\right]$ (23)
$=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{mn^{\prime}M}{mn^{\prime}N}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$ (24)
$=E_{n^{\prime}}\left[\frac{M}{N}\sum_{{m=1}}^{min(M,n^{\prime})}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$ (25)
$=M/N=\text{Recall}@K.$ (26)
A.5 Compared Methods
We utilize two representative recommendation models, Matrix Factorization (MF) (Koren, Bell, and Volinsky 2009) and graph neural network model LightGCN (He et al. 2020), as the backbones. Based on the Theorem 2, We devise NegS, which optimizes the backbone model through BCE loss accompanied by the hard negative sampling, and compare NegS with following debiasing methods:
  • IPS (Schnabel et al. 2016) weights the loss of each sample with an inversed propensity. We follow (Chen et al. 2021) and calculate the propensity with naive bayes.
  • DR (Wang et al. 2019) integrates imputed errors and propensities while mitigating the impact of high propensity variance.
  • AutoDebias (Chen et al. 2021) propose a meta-learning-based algorithm to learn debiasing parameters from a small set of randomly-exposed data.
A.6 Metrics
We use Recall metric to evaluate models for our following experiments. The Recall metric is the most important metric in the candidate generation stage of industrial recommendation systems. It is also a commonly used metric in research on debiasing recommendations. In future work, we will investigate the unbiased estimation of other Top-$K$ metrics, like Precision and NDCG on the fully-exposed dataset to further extend our conclusions.
A.7 Setup
All the models are optimized by Adam optimizer. We use grid search to find the best hyperparameters. The search space of learning rate and weight decay are [1e-4, 1e-3, 1e-2, 1e-1]. For the proposed NegS, we tune $H/U$ and $U$ within the ranges of [0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1] and [50,100,200] respectively. For details on the dataset, see Appendix A.1. For details on the baseline, see Appendix A.5. For details on the matrics, see Appendix A.6.

### v2

@@
-Debias Can be Unreliable: Mitigating Bias Issue in Evaluating Debiasing Recommendation
-Abstract
+Debias Can Be Unreliable: Mitigating Bias in Evaluating Debiasing Recommendation
+Abstract.
 Recent work has improved recommendation models remarkably by equipping them with debiasing methods.
@@
 However, in this study, we reveal that traditional evaluation scheme is not suitable for randomly-exposed datasets, leading to inconsistency between the Recall performance obtained using randomly-exposed datasets and that obtained using fully-exposed datasets.
-Such inconsistency indicates the potential unreliability of experiment conclusions on previous debiasing techniques and calls for unbiased Recall evaluation using randomly-exposed datasets.
 To bridge the gap, we propose the Unbiased Recall Evaluation (URE) scheme, which adjusts the utilization of randomly-exposed datasets to unbiasedly estimate the true Recall performance on fully-exposed datasets.
 We provide theoretical evidence to demonstrate the rationality of URE and perform extensive experiments on real-world datasets to validate its soundness.
-1 Introduction
-Recommendation models suffer from various biases such as selection bias (Marlin et al.
-2007; Hemminki and Sihvo 1993; Huang, Oosterhuis, and de Rijke 2022; Vernade and Cappé 2015) and popularity bias (Abdollahpouri and Mansoury 2020; Zhang et al.
-2021; Zhu et al.
-2021; Abdollahpouri et al.
-2021), leading to deviations from true user preferences even harmful impacts to Web services (Chen et al.
-2023; Klimashevskaia et al.
-2023; Klimashevskaia 2022; Kowald, Schedl, and Lex 2020), such as amplifying the long-tail effect and hurting user trust (Abdollahpouri, Burke, and Mobasher 2017; Abdollahpouri et al.
-2019; Abdollahpouri 2019).
+Our code and datasets are available at https://anonymous.4open.science/r/URE-6B64/.
+Debiasing Recommendation; Evaluation
+journalyear: 2025copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715548isbn: 979-8-4007-1331-6/2025/04ccs: Information systems Recommender systems
+1. Introduction
+Recommendation models suffer from various biases leading to deviations from true user preferences even harmful impacts to Web services (Chen et al., 2023; Zhang et al., 2021; Schnabel et al., 2016; Li et al., 2023; Pan et al., 2023).
 Undoubtedly, user’s true preferences can be accurately measured by the fully-exposed dataset which includes users’ feedback (e.g., ratings) across all items.
-The gold standard for evaluating debiasing techniques is thus utilizing such a fully-exposed dataset to assess the performance of recommendations (Gao et al.
-2022a).
-However, due to the unavailability of fully-exposed datasets in practice (Gao et al.
-2022b), randomly-exposed datasets, which contain users’ feedback on some randomly selected items, are widely employed for debiasing evaluation (Liu et al.
-2020; Yuan et al.
-2019; Rosenfeld, Mansour, and Yom-Tov 2017; Bonner and Vasile 2018; Schnabel et al.
-2016; Wang et al.
-2019; Ai et al.
-2018; Chen et al.
-2021).
-Notably, the amount of feedback in randomly-exposed datasets is also extremely limited, because randomly exposing items to collect user feedback is highly expensive, and the collection process can severely damage a platform’s revenue.
-Figure 1: Preliminary analysis on KuaiRec, an only existing fully-exposed dataset, validating the inconsistency between Recall@$\overline{K}$ on the randomly-exposed data ($l$ = 80 samples per user) and Recall@$K$ on the fully-exposed data ($L$ = 3327 samples per user) when $K$ is small.
-We rank four AutoDebias models with different hyperparameters w.r.t.
-Recall@$\overline{5}$ and Recall@$K$, respectively.
-Notably, $\overline{K}\in$ [1,2,…$l$], $K\in$ [1,2,…$L$], where $l<<L$.
+The gold standard for evaluating debiasing techniques is thus utilizing such a fully-exposed dataset to assess the performance of recommendations (Gao et al., 2022a).
+However, due to the unavailability of fully-exposed datasets in practice (Gao et al., 2022b), randomly-exposed datasets, which contain users’ feedback on some randomly selected items, are widely employed for debiasing evaluation (Schnabel et al., 2016; Liu et al., 2020; Rosenfeld et al., 2017).
+Notably, the amount of feedback in randomly-exposed datasets is extremely limited, because randomly exposing items to collect user feedback is highly expensive, and the collection process can severely damage a platform’s revenue.
 Previous work calculate widely-used Recall metric on randomly-exposed datasets in the same way as on fully-exposed datasets, which we refer to as the traditional evaluation scheme in subsequent descriptions.
-Nevertheless, we claim that traditional evaluation scheme is not suitable for randomly-exposed datasets when evaluating debiasing recommendation.
-Specifically, leveraging the only existing full-exposed dataset, KuaiRec (Gao et al.
-2022b), our preliminary analysis reveals significant disparities between the evaluation results obtained from the randomly-exposed dataset and those derived from the fully-exposed dataset.
-Figure 1 is a quick example that presents the rankings of different models (i.e., AutoDebias models with different hyperparameters) based on the Recall@$\overline{5}$ on the randomly-exposed dataset and Recall@$K$ on the fully-exposed dataset.
-As can be seen, the model ranking deduced from Recall@$\overline{5}$ is different from those derived from Recall@$K$ when $K\leq 100$.
-Given that user typically browse only a limited number of items during an online deployment stage, Recall@$K$ with a small value of $K$ (e.g., $K=50$) on the fully-exposed dataset represents desired debiasing quality and correct model ranking.
-Consequently, the conclusion derived from the traditional evaluation scheme on the randomly-exposed dataset such as Recall@$\overline{5}$ might be unconvincing.
-To comprehensively investigate the underlying factors contributing to the inconsistency between the Recall@$\overline{K}$ on the randomly-exposed dataset and Recall@$K$ on the fully-exposed dataset, we undertake a theoretical analysis (cf.
+Nevertheless, we claim that traditional evaluation scheme is not suitable for computing Recall metric on randomly-exposed datasets when evaluating debiasing recommendation.
+This is due to the significant disparities between the Recall obtained from the randomly-exposed dataset and those derived from the fully-exposed dataset.
+To comprehensively investigate the underlying factors contributing to the inconsistency, we undertake a theoretical analysis (cf.
 Theorem 2) of the evaluation performed on a randomly-exposed dataset.
 Our analysis demonstrates that the Recall@$\overline{K}$ on the randomly-exposed dataset (e.g., Recall@$\overline{5}$) has a weak correlation with the Recall@$K$ using a small value of $K$ on the fully-exposed dataset (e.g., Recall@50).
-Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al.
-2023).
-In addition, we conduct extensive experiments on real-world datasets to further validate the variance of such correlation (cf.
+Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al., 2023).
+In addition, we conduct extensive experiments on real-world datasets to validate such correlation (cf.
 Figure 1).
-Given that the traditional evaluation scheme on the randomly-exposed dataset has limited capacity to reflect the desired Recall@$K$ performance on the fully-exposed dataset, it is inadequate to conclude whether previous debiasing methods have effectively mitigated biases.
-The observed performance gain of debiasing methods on the traditional evaluation scheme can be attributed to optimizing Recall@$K$ under a large value of $K$ on the fully-exposed dataset using techniques that lack genuine debiasing effects.
-For instance, hard negative sampling that theoretically optimizes Recall@$K$ with a large $K$ on fully-exposed datasets (Shi et al.
-2023), can outperform the representative debiasing method (Chen et al.
-2021) under the traditional evaluation scheme (see Table 3).
-Considering the cost of constructing fully-exposed dataset, unbiasedly estimating the Recall@$K$ performance with a small value of $K$ on the fully-exposed dataset from randomly-exposed datasets is crucial for the evaluation of debiasing methods.
-Figure 2: Illustration of URE.
-In randomly-exposed data, $n$ is the number of positive items with existing feedback and $m$ is the number of positive items with existing feedback ranked within top-$K$ of all items.
-Towards this goal, we propose the Unbiased Recall Evaluation (URE) scheme as illustrated in Figure 2.
-We set the target as estimating the Recall@$K$ on fully-exposed data, i.e., the ratio of items with positive feedback ranked before position $K$, denoted as $\frac{M}{N}$.
+Thus, the conclusion derived from the traditional evaluation scheme on the randomly-exposed dataset such as Recall@$\overline{5}$ might be unconvincing as it is inadequate to conclude whether previous debiasing methods have effectively mitigated biases.
+Towards this goal, we propose the Unbiased Recall Evaluation (URE) scheme.
+We set the target as estimating the Recall@$K$ on fully-exposed data, i.e., the ratio of items with positive feedback ranked before position $K$.
 Considering that a randomly-exposed dataset only includes user feedback on randomly selected items, we can treat it as a representative subset of the fully-exposed dataset, wherein both positive and negative feedback are missing at random.
 In URE, we sort the model’s prediction scores of all candidate items and utilize the score of $(K+1)$-th item as the threshold to calculate the positive ratio on the randomly-exposed dataset, where the positive ratio is the proportion of items with positive feedback in the randomly-exposed dataset that rank before the threshold.
-Hence such a positive ratio can be used to unbiasedly estimate $\frac{M}{N}$ on fully-exposed dataset.
+Hence averaging such a positive ratio across all users can be used to unbiasedly estimate Recall@$K$ on a fully-exposed dataset.
 This is due to the models’ prediction information of all candidate items, which is not considered by traditional evaluation scheme.
 Through theoretical analysis and extensive experiments, we validate that URE can unbiasedly estimate the Recall@$K$ on fully-exposed data with a desired value of $K$.
-In this work, our contributions are threefold.
+The main contributions are summarized as follows:
   • We reveal the inconsistency between the widely adopted traditional evaluation scheme on the randomly-exposed dataset and fully-exposed dataset (the gold standard evaluation), questioning empirical conclusions on existing debiasing methods in previous studies.
-  • We theoretically and empirically analyze the relation between the Recall@$\overline{K}$ on the randomly-exposed dataset and the Recall@$K$ on the fully-exposed dataset.
-  • We propose the URE scheme to unbiasedly estimate the target Recall@$K$ performance with randomly-exposed data, and conduct extensive theoretical and empirical analysis, validating the effectiveness of the URE scheme.
-2 Preliminary
-Table 1: Notations and definitions.
-Notations Annotations
-$D_{full}$ A fully-exposed dataset
-$D_{rand}$ A randomly-exposed dataset
-$L/l$ The amount of feedback from a single user
-$K$ Top-K on $D_{full}$, $K\in$ [0,1,…$L$]
-$\overline{K}$ Top-K on $D_{rand}$, $\overline{K}\in$ [0,1,…$l$] ($l<<L$)
-Recall@$K$ Recall metric at position $K$ on $D_{full}$
-Recall@$\overline{K}$ Recall metric at position $\overline{K}$ on $D_{rand}$
-$\widehat{\text{Recall@}K}$ The estimation of $\text{Recall}@K$ by URE scheme
-Given a user ($u$) set and an item ($i$) set, fully-exposed dataset $D_{full}$ includes the user feedback $r_{u,i}$ (e.g., ratings and clicks) on all candidate items, which is usually unavailable in practice.
-The only existing $D_{full}$ for research is KuaiRec (Gao et al.
-2022b).
-All candidate items can be categorized into items with positive feedback that user likes ($r_{u,i}=1$) and items with negative feedback ($r_{u,i}=0$) based on user’s true interests.
-For the evaluation of debiasing recommendation, researchers mostly resort to randomly-exposed datasets $D_{rand}$ as a proxy to evaluate debiasing methods.
-However, randomly-exposed dataset $D_{rand}$ collects the user feedback on some randomly selected items ($\in$ all candidate items) through a costly randomized intervention experiment (Liu et al.
-2020; Yuan et al.
-2019; Rosenfeld, Mansour, and Yom-Tov 2017; Bonner and Vasile 2018; Schnabel et al.
-2016; Wang et al.
-2019; Ai et al.
-2018; Chen et al.
-2021), i.e., a large amount of user feedback is ramdomly missing.
-During the evaluation stage, recommendation models ($\hat{y}=M(u,i)$) rank the items with feedback based on $\hat{y}$ to generate a personalized ranking list $Rank_{u}$ for each user.
-In traditional evaluation scheme on $D_{full}$, we can obtain the true performance of debiased recommendation models by calculating the Recall@$K$.
-Notably, The Recall metric is the most important metric in the candidate generation stage of industrial recommendation systems.
-It is also a commonly used metric in research on debiasing recommendations.
-Recall@$K$ on $D_{full}$ is obtained by the equation 1.
-$\text{Recall}@K=\frac{|{k\in Rank_{u}+:k\leq K}|}{|Rank_{u}+|},K\in[1,2,...,L],$ (1)
-where the $Rank_{u}+$ collects the positions of all items with positive feedback within $Rank_{u}$.
-Unfortunately, as mentioned above, fully-exposed dataset is unavailable on most recommendation platforms.
-The Recall@$K$ in the equation 1 is not supported if the feedback of any candidate item is missing.
-For the evaluation of debiasing recommendation, researchers employ traditional evaluation scheme on $D_{rand}$ to calculate the Recall performance, denoted as Recall@$\overline{K}$.
-Figure 3: The Pearson correlation coefficients or Kendall coefficients between Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$.
-We highlight the value of $K$ when each curve reaches its maximum correlation coefficient as $K_{max}$.
-(a) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5.
-(b) The effect of $\overline{K}$, where $\overline{N}$ is fixed to 80.
-(c) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5.
-The correlations are statistically significant with $p$ $<$ 0.01.
-$\begin{split}\text{Recall}@\overline{K}=\frac{|{k\in\overline{Rank_{u}}+:k\leq\overline{K}}|}{|\overline{Rank_{u}}+|},\\ \overline{K}\in[1,2,...,l],\quad l<<L,\end{split}$ (2)
-where $\overline{Rank_{u}}+$ contains the positions of items with positive feedback of user $u$ in $\overline{Rank_{u}}$.
-However, in this work, we find that Recall@$\overline{K}$ on the randomly-exposed dataset is inconsistent with Recall@$K$ on fully-exposed dataset.
-Therefore traditional evaluation scheme is not suitable for randomly-exposed datasets.
-We will detail the inconsistency and unreliability of such a traditional evaluation scheme in the following sections.
+  • We propose the URE, a novel scheme to unbiasedly estimate the target Recall@$K$ performance with randomly-exposed data.
+  • We conduct extensive theoretical and empirical analysis, validating the effectiveness of the URE scheme.
+2. Preliminary
+Given a fully-exposed dataset (denoted as $D_{full}$), we can determine the true performance of debiasing models by calculating the Recall formula, denoted as denoted as Recall@$K$:
+(1) $\begin{split}\text{Recall}@K=\frac{|{k\in Rank_{u}+:k\leq K}|}{|Rank_{u}+|},K\in[1,2,...,L],\end{split}$
+where the $Rank_{u}+$ compiles the positions of all items with positive feedback within the personalized ranking list recommended by the models.
+However, The Recall@$K$ is not supported due to the unavailability of $D_{full}$ (The only existing $D_{full}$ for research is KuaiRec (Gao et al., 2022b)).
+In this case, researchers employ the same formula to compute Recall on a randomly-exposed dataset (denoted as $D_{rand}$).
+The resulting Recall is referred to as Recall@$\overline{K}$, where $\overline{K}\in[1,2,...,l],l<<L$.
+This evaluation approach for debiasing models is termed the ”traditional evaluation scheme on $D_{rand}$”.
 […]
-In this section, we first conduct theoretical analysis to discover why the traditional evaluation scheme on randomly-exposed data is unreliable by inspecting the relationship between Recall@$K$ metric on fully-exposed data and Recall@$\overline{K}$ metric on randomly-exposed data.
-Then, we substantiate their correlation through empirical validation.
-3.1 Theoretical Guarantee
-In this subsection, we give the theoretical derivation of the relationship between the Recall@$K$ on $D_{full}$ and the Recall@$\overline{K}$ on $D_{rand}$.
+In this section, we conduct theoretical analysis and empirical validation to discover why the traditional evaluation scheme on $D_{rand}$ is unreliable by inspecting the relationship between Recall@$K$ on $D_{full}$ and Recall@$\overline{K}$ on $D_{rand}$.
+3.1. Theoretical Guarantee
+In this subsection, we give the theoretical derivation of the relationship between Recall@$K$ on $D_{full}$ and Recall@$\overline{K}$ on $D_{rand}$.
 Theorem 1.
@@
 Then, when $K$ on $D_{full}$ and $\overline{K}$ on $D_{rand}$ satisfy $\overline{K}=\frac{\overline{N}}{N}\cdot K$, we have:
-$E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]-\text{Recall@}K\right]=0,$ (3)
+(2) $E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]-\text{Recall@}K\right]=0,$
 where $E$ denotes the Expectation funtion.
-The proof of Theorem 3 can be found in Appendix A.2.
+Proof.
+We first calculate the $E_{\mathcal{R}}[\text{Recall@}K]$.
+Since we consider all possible ranking, the possibility of the x positive samples contained in the top K is $\frac{C_{N^{+}}^{x}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$ where $C^{n}_{N}:=\frac{N!}{(N-n)!n!}$ is the number of combinations obtained by randomly sampling $n$ samples from $N$.
+Thus, we have:
+(3) $E_{\mathcal{R}}\left[\text{Recall}@K\right]=\sum_{x=1}^{min(K,N^{+})}\frac{x}{N^{+}}\frac{C^{x}_{N^{+}}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$
+(4) $=\frac{K}{N}\left[\sum_{x=1}^{min(K,N^{+})}\left[C_{N^{+}-1}^{x-1}C_{(N-1)-(N^{+}-1)}^{(K-1)-(x-1)}/C^{K-1}_{N-1}\right]\right]$
+(5) $=\frac{K}{N}\left[\ C_{N-1}^{K-1}/C_{N-1}^{K-1}\right]=\frac{K}{N}.$
+As for the $E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]$, we have:
+(6) $E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]=E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right].$
+Since we consider all possible permutations, $E_{\mathcal{R},\mathcal{S}_{\overline{N}}}\left[\text{Recall@}\overline{K}|{N^{+}}^{\prime}\right]$ can be transformed into the scenario of randomly selecting $\overline{K}$ samples from $\overline{N}$ samples ($N^{+}$ positive samples among these $N$ samples) and determining how many of the selected $\overline{K}$ samples are positive.
+The problem is similar to Eq 3.
+Thus we have:
+(7) $E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right]$
+(8) $=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x}{{N^{+}}^{\prime}}\frac{C^{x}_{{N^{+}}^{\prime}}C^{\overline{K}-x}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}}^{\overline{K}}}\right]$
+(9) $=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\frac{\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}-1}^{\overline{K}-1}}\right]$
+(10) $=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\right]=\frac{\overline{K}}{\overline{N}}=\frac{K}{N}=E_{\mathcal{R}}[\text{Recall@}K].$
 From Theorem 2, we can observe that the Recall@$\overline{K}$ on $D_{rand}$ will exhibit a strong correlation with the Recall@$K$ on $D_{full}$ if the value of $K$ is around $\frac{{N}}{\overline{N}}\cdot\overline{K}$.
-It should be noted that $\overline{N}$ is much smaller than $N$ since the randomly-exposed dataset is a sparsely sampled subset of the fully-exposed dataset.
+It should be noted that $\overline{N}$ is much smaller than $N$ since the $D_{rand}$ is a sparsely sampled subset of $D_{full}$.
 Therefore, given a specific value of $\overline{K}$, the Recall@$\overline{K}$ on $D_{rand}$ only exhibits a strong correlation with the Recall@$K$ with much larger $K$ values on $D_{full}$.
-However, given that user typically browse only a limited number of items during an online deployment stage, Recall@$K$ with a small value of $K$ (e.g., $K=50$) on the $D_{full}$ represents desired debiasing quality and correct model ranking.
-Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al.
-2023).
-3.2 Empirical Results
-In this subsection, we conduct empirical experiments to validate the relationship between Recall@$K$ metric on fully-exposed data and Recall@$\overline{K}$ metric on randomly-exposed data.
-We train multiple AutoDebias (Chen et al.
-2021) models w.r.t.
+However, given that user typically browse only a limited number of items during an online deployment stage, the performance cutting at a small value of $K$ is more important (Chen et al., 2023).
+Figure 1.
+The correlation coefficients between Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$.
+(a) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5.
+(b) The effect of $\overline{K}$, where $\overline{N}$ is fixed to 80.
+3.2. Empirical Results
+In this subsection, we conduct empirical experiments to validate the relationship between Recall@$K$ on $D_{full}$ and Recall@$\overline{K}$ on $D_{rand}$.
+We train multiple AutoDebias (Chen et al., 2021) models w.r.t.
 different hyperparameters and evaluate their Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$.
-Then we calculate Pearson and Kendall coefficients between Recall@$\overline{K}$ and Recall@$K$ to measure their linear correlation and ranking correlation (Abdi 2007), respectively.
+Then we calculate correlation coefficients between Recall@$\overline{K}$ and Recall@$K$ to measure their correlation.
 The results are shown in Figure 1.
@@
   • When fixing $\overline{K}$, the value of $K_{max}$ decreases as the sample number $\overline{N}$ increases.
-As shown in Figure 3a and Figure 3c, the inverse relationship between $K_{max}$ and $\overline{N}$ is consistent with the conclusion in Eq.
+As shown in Figure 1a, the inverse relationship between $K_{max}$ and $\overline{N}$ is consistent with the conclusion in Eq.
 (2) that $K=\frac{N}{\overline{N}}\cdot\overline{K}$.
@@
 However, as we discussed before, due to the frequent occurrence of $\overline{N}\ll N$, even when considering Recall@$\overline{K}$ with $\overline{K}=1$, it is still not possible to unbiasedly and accurately estimate Recall@$K$ for small values of $K$.
-  • When fixing $\overline{K}$, the highest Kendall correlation coefficient increase as sample number $\overline{N}$ increase.
-Figure 3c shows that when sample number $\overline{N}$ is small, Kendall correlation coefficient will be small for any Recall@$K$.
-Low Kendall correlation coefficient means that if model A has a higher Recall@$\overline{K}$ value on $D_{rand}$ than model B, it does not necessarily indicate that A performs better than B with Recall@$K$ on $D_{full}$.
-Therefore, it is unreliable to use the Recall@$\overline{K}$ on $D_{rand}$ for model selection.
-In conclusion, we prove from both theoretical and experimental aspects that there is a relationship between Recall@$K$ on the fully-exposed dataset and Recall@$\overline{K}$ on the randomly-exposed dataset.
-Specifically, Recall@$\overline{K}$ is associated with Recall@$K$ when the value of $K$ is relatively large.
-Given that the gold standard for debiasing recommendation desires Recall@$K$ with small values of $K$ (e.g., $K=50$), using traditional evaluation scheme on randomly-exposed dataset directly is problematic.
-4 Entrapment Study
-[…]
-4.1 Hard Negative Sampling
-In this subsection, we begin by introducing the concept of hard negative sampling.
-Subsequently, we present a theorem delineating the relationship between hard negative sampling and the optimization of Recall@$K$ on $D_{full}$ with the desired value of $K$.
-Hard negative sampling refers to oversampling high-scored negative items for training.
-Taking the representative method DNS($H$, $U$) (Shi et al.
-2023) as an example, it first uniformly selects $U$ negative samples and then selects the top $H$ hard samples with the highest prediction scores by the model for training.
-Particularly, as $H/U$ decreases, it implies that the selected hard negative samples become more challenging for the model.
-We thus define $H/U$ as the hardness of hard negative sampling.
-We can present a theorem elucidating the relationship between hard negative sampling and optimizing Recall@$K$ on $D_{full}$.
-Theorem 2.
-When the error function $\delta$ is the BCE loss function, the model equipped with hard negative sampling DNS($H$, $U$) approximately optimizes Recall@$K$ metric.
-Meanwhile, the larger $H/U$ is, the Recall@$K$ metric with a larger $K$ on $D_{full}$ is optimized.
-The proof can be found in Appendix A.3.
-Based on the Theorem 2, we conclude that the hard negative sampling approach could optimize Recall@$K$ with large values of $K$ on $D_{full}$ by using large $H/U$.
-Table 2: The statistics of datasets.
-Dataset ##\##Users ##\##Items Biased train test
-Yahoo!R3 15,400 1,000 311,704 48,600 ($D_{rand}$)
-KuaiRec 1411 3327 1,934,404 2,742,166 ($D_{full}$)
-Table 3: The comparison between NegS and representative methods with debiasing techniques under the traditional evaluation scheme on $D_{rand}$.
-The best results are in bold and the second best are underlined.
-The results are statistically significant with $p<$ 0.01.
-recall@$\overline{5}$ MF LightGCN
-Yahoo!R3 KuaiRec Yahoo!R3 KuaiRec
-Bias 0.7085 0.3546 0.7883 0.4859
-IPS 0.6572 0.3102 0.7584 0.4889
-DR 0.7195 0.5465 0.7576 0.4849
-AutoDebias 0.8085 0.5482 0.8046 0.5181
-NegS 0.8330 0.5358 0.8141 0.5288
-[…]
-[…]
-4.2.1 Experimental Setup.
-To assess the reliability of the traditional evaluation scheme on $D_{rand}$, we use two widely-used datasets for our following experiments: Yahoo!R3 and KuaiRec.
-The statistics related to these datasets can be found in Table 2.
-The KuaiRec is the only existing full-exposed dataset that also can be used to construct randomly-exposed set and training set.
-However, Yahoo!R3 only consists of a biased set for training and an randomly-exposed set for testing.
-We utilize two representative recommendation models, Matrix Factorization (MF) (Koren, Bell, and Volinsky 2009) and graph neural network model LightGCN (He et al.
-2020), as the backbones.
-We compare NegS with following debiasing methods: IPS (Schnabel et al.
-2016), DR (Wang et al.
-2019), AutoDebias (Chen et al.
-2021).
-All the models are optimized by Adam optimizer.
-AutoDebias uses random search to find the best hyperparameters while other models use grid search.
-The search space of learning rate and weight decay are [1e-4, 1e-3, 1e-2, 1e-1].
-For the proposed NegS, we tune $H/U$ and $U$ within the ranges of [0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1] and [50,100,200] respectively.
-The detialed setup can be found in Appendix A.7.
-4.2.2 Observations.
-We compare the NegS to debiasing baselines using the traditional evaluation scheme on $D_{rand}$ in Table 3.
-For NegS, we utilize a large $H/U$ to obtain the best results.
-As shown in Table 3, we highlight that the NegS exhibits comparable or even better performance compared to the debiasing models on Recall@$\overline{K}$ metric on two datasets.
-From this example, we can infer that a model that optimizes the Recall@$K$ metric with a large $K$ on $D_{full}$, even without any debiasing technologies, can also achieve remarkable performance under the traditional evaluation scheme on $D_{rand}$.
-This shows that the traditional evaluation scheme on $D_{rand}$ cannot accurately identify effective debiasing models.
-As a result, the conclusions drawn from evaluating debiasing techniques in this manner may be unreliable, necessitating significant effort in future work for further investigation.
-5 URE scheme
-Figure 4: Illustration of the proposed URE scheme on unbiasedly estimating Recall@$K$ on $D_{full}$.
-Considering the impractical accessibility of $D_{full}$, the pivotal role of $D_{rand}$ emerges in enabling an unbiased estimation of Recall@$K$ on $D_{full}$.
-Nevertheless, as mentioned above, our contention is that the traditional evaluation scheme on $D_{rand}$ exhibits bias.
-In response to this challenge, we introduce a pioneering evaluation scheme coined URE, delineated in Figure 4.
+In conclusion, Recall@$\overline{K}$ is associated with Recall@$K$ when the value of $K$ is relatively large.
+Given that the gold standard for debiasing recommendation desires Recall@$K$ with small values of $K$ (e.g., $K=50$), Recall@$\overline{K}$ is problematic.
+4. URE scheme
+In this section, we introduce a pioneering evaluation scheme coined URE, delineated in Figure 2.
 URE relies solely on $D_{rand}$ to achieve an unbiased estimation of Recall@$K$ on $D_{full}$, which consists of the following steps:
@@
   • Locate the ($K$+1)-th item in Sorted ItemID and its corresponding prediction $\hat{y}$.
-  • Compute the $\widehat{\text{Recall@}K}$ for this user:
-    $\widehat{\text{Recall@}K}=\frac{m}{n},$
-    where $n$ denotes the number of items with positive label in Sorted $D_{rand}$, and $m$ denotes the number of these items with higher $\hat{y}$ than the ($K$+1)-th item.
+  • Compute the $\widehat{\text{Recall@}K}$ for this user: $\widehat{\text{Recall@}K}=\frac{m}{n},$ where $n$ denotes the number of items with positive label in Sorted $D_{rand}$, and $m$ denotes the number of these items with higher $\hat{y}$ than the ($K$+1)-th item.
   • We report the averaged $\widehat{\text{Recall@}K}$ across all users as an unbiased estimate of the model’s overall Recall@$K$ performance on $D_{full}$.
@@
 (URE Unbiasness) For all given $K$, we have:
-$E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall@}K}\right]=\text{Recall}@K,$ (4)
+(11) $E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall@}K}\right]=\text{Recall}@K,$
 where $\mathcal{S}_{\overline{N}}$ denotes the set of all possible random samplings of size $\overline{N}$.
-The proof can be found in Appendix A.4. This signifies that the estimates derived from URE are unbiased.
-In addition to the theoretical analysis, we provide an intuitive explanation to show how our URE can unbiasedly estimate the Recall@$K$ on $D_{full}$.
-The core reliance is that $D_{rand}$ represents a random selection from the sorted list of $D_{full}$.
-Hence, the ratio of items with positive label in $D_{rand}$ which have a higher $\hat{y}$ than the ($K$+1)-th item can be used to unbiasedly estimate the positive ratio of $D_{full}$.
-Our proposed URE scheme, despite its apparent simplicity, can generate outstanding predictive accuracy.
-This is due to the models’ prediction information of all candidate items, which is not considered by traditional evaluation scheme.
-We will demonstrate the effectiveness of this scheme in following sections.
-Figure 5: (a) The Kendall rank coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$.
-(b) The Kendall rank coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$.
-We mark $K_{max}$ when every curve reaches its highest coefficient.
-The rank correlations in Figure (a) and (b) are statistically significant with $p<$ 0.01.
-6 Experiments
-Table 4: The evaluation results by using the URE scheme.
-We also show the Recall@$\overline{5}$ under the traditional evaluation scheme for comparison.
-The best results are in bold and the second best are underlined.
-Notably, the Recall@$5$ metric is not available in Yahoo!R3, as it lacks a fully-exposed dataset.
-The results are statistically significant with $p<$ 0.01.
-Backbone Model KuaiRec Yahoo!R3
-Traditional scheme Fully-Exposed Dataset URE scheme Fully-Exposed Dataset Traditional scheme URE scheme
-Recall@$\overline{5}$ Recall@$5$ $\widehat{\text{Recall@5}}$ $\widehat{\text{Recall@500}}$ Recall@500 Recall@$\overline{5}$ $\widehat{\text{Recall@5}}$ $\widehat{\text{Recall@500}}$
-MF Bias 0.3546 0.0215 0.0213 0.3280 0.3237 0.7085 0.0131 0.6777
-IPS 0.3102 0.0213 0.0230 0.2771 0.2743 0.6572 0.0095 0.6236
-DR 0.5465 0.0353 0.0362 0.5077 0.5050 0.7195 0.0075 0.6879
-AutoDebias 0.5482 0.0355 0.0352 0.5129 0.5097 0.8085 0.0550 0.7849
-NegS 0.5358 0.0327 0.0326 0.5052 0.5012 0.8330 0.0490 0.8001
-LightGCN Bias 0.4859 0.0257 0.0267 0.4503 0.4482 0.7883 0.0491 0.7680
-IPS 0.4889 0.0266 0.0273 0.4539 0.4522 0.7584 0.0457 0.7367
-DR 0.4849 0.0262 0.0267 0.4500 0.4457 0.7576 0.0455 0.7373
-AutoDebias 0.5181 0.0317 0.0323 0.4571 0.4558 0.8046 0.0592 0.7820
-NegS 0.5288 0.0237 0.0245 0.4954 0.4944 0.8141 0.0616 0.7913
-In this section, we conduct experiments to verify the effectiveness of the URE scheme and re-evaluate the debiasing methods using the proposed URE scheme.
-6.1 Unbiasedness Analysis
-In this subsection, we test the AutoDebias model using the KuaiRec dataset to verify that the proposed $\widehat{\text{Recall@}K}$ in the URE scheme is an unbiased estimation of Recall@$K$ on $D_{full}$.
-In Figure 5a, we investigate whether there is order preservation between Recall@$K$ and our $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$.
-Concretely, we randomly select model’s hyperparameter combinations to train the AutoDebias model on the Kuairec dataset.
-Then we calculate the Kendall rank correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$, as well as Recall@$\overline{K}$ to demonstrate the ranking consistency.
+This signifies that the estimates derived from URE are unbiased.
+Proof.
+Since Recall is only related to positive samples, we can only consider positive samples.
+Let’s assume that in $D_{full}$, there are a total of $N$ positive samples, out of which $M$ samples are ranked higher than the sample ranked at $K+1$ in Sorted ItemID in Figure 2.
+Assuming that we have randomly sampled $\overline{N}$ samples, which have $n$ positive samples and the rest are negatives.
+Since the sampling process is random, the probability of the number $m$ of positive samples that rank higher than the sample ranked at $K+1$ in Sorted ItemID can be calculated as $(C^{m}_{M}C^{n-m}_{N-M}/C^{n}_{N})$, thus we have:
+Figure 2. Illustration of the proposed URE scheme on unbiasedly estimating Recall@$K$ on $D_{full}$.
+(12) $E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall}@K}\right]$ $=E_{n^{\prime}}\left[E_{\mathcal{S}_{\overline{N}}}[m/n|n^{\prime}]\right]$
+(13) $=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{mn^{\prime}M}{mn^{\prime}N}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$
+(14) $=E_{n^{\prime}}\left[\frac{M}{N}\sum_{{m=1}}^{min(M,n^{\prime})}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$
+(15) $=M/N=\text{Recall}@K.$
+Figure 3.
+(a) The correlation coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$.
+(b) The correlation coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$.
+5. Experiments
+Unbiasedness Analysis.
+In this subsection, we test the AutoDebias model using the KuaiRec dataset to verify that the $\widehat{\text{Recall@}K}$ in the URE scheme is an unbiased estimation of Recall@$K$ on $D_{full}$.
+In Figure 3a, we calculate the correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$, as well as Recall@$\overline{K}$ to demonstrate the ranking consistency.
 As shown in Figure 3a, when $K$ is small (e.g.
 $K=30$), both Recall@$\overline{1}$ and Recall@$\overline{5}$ exhibits poor correlations with Recall@$K$.
-In contrast, the Kendall correlation coefficient between the estimated $\widehat{\text{Recall@}K}$ and Recall@$K$ is higher than 0.9.
+In contrast, the correlation coefficient between the estimated $\widehat{\text{Recall@}K}$ and Recall@$K$ is higher than 0.9.
 This means that the model we choose based on $\widehat{\text{Recall@}K}$ is more likely to perform better on Recall@$K$ on $D_{full}$, which further demonstrates the effectiveness of the URE scheme.
-In Figure 5b, we report the Kendall rank correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$.
+In Figure 3b, we report the correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$.
 As we desired, only when K is the same, Recall@$K$ and $\widehat{\text{Recall@}K}$ have the highest correlation coefficients.
 That is, URE allows us to select the best debiased model based on the Recall@$K$ with a desired $K$.
-Based on our discussion, we show that our URE scheme unbiasedly estimates Recall@$K$ for different values of $K$.
-Moreover, the high Kendall coefficients between $\widehat{\text{Recall@}K}$ and Recall@$K$ ensure an accurate debiased model selection.
-6.2 Re-evaluation Results
-In this subsection, we re-evaluate some classical debiasing methods, and report the results in Table 4. We have the following key findings:
-  • Based on the results on KuaiRec dataset, we observe that the evaluation result $\widehat{\text{Recall@5}}$ of the URE scheme aligns closely with the true Recall@5 on the fully-exposed dataset.
+Based on our discussion, we show that our URE scheme unbiasedly estimates Recall@$K$ for different values of $K$ and ensure an accurate debiased model selection.
+Table 1. The evaluation results by URE. Notably, the Recall@$5$ is not available in Yahoo!R3, as it lacks a fully-exposed dataset. The results are statistically significant with $p<$ 0.01.
+Model KuaiRec Yahoo!R3
+Traditional scheme Fully-Exposed Dataset URE scheme Traditional scheme URE scheme
+Recall@$\overline{5}$ Recall@$5$ $\widehat{\text{Recall@5}}$ Recall@$\overline{5}$ $\widehat{\text{Recall@5}}$
+MF 0.4859 0.0257 0.0267 0.7085 0.0131
+IPS 0.4889 0.0266 0.0273 0.6572 0.0095
+DR 0.4849 0.0262 0.0267 0.7195 0.0075
+AutoDebias 0.5181 0.0317 0.0323 0.8085 0.0550
+DNS 0.5288 0.0237 0.0245 0.8330 0.0490
+Re-evaluation Results.
+In this subsection, we re-evaluate some classical debiasing methods under backbone MF (Koren et al., 2009), including IPS (Schnabel et al., 2016), DR (Wang et al., 2019), AutoDebias (Chen et al., 2021), DNS (Shi et al., 2023) on two widely-used datasets (cf.
+Table 2) in the debiasing research domain, and report the results in Table 1.
+We have the following key findings: (1) Based on the results on KuaiRec, we observe that the evaluation result $\widehat{\text{Recall@5}}$ of the URE scheme aligns closely with the true Recall@5 on $D_{full}$.
 This observation suggests a strong unbiasedness in the URE scheme.
-In contrast, the evaluation results of the traditional scheme exhibit a significant discrepancy from the true Recall@5, and it fails to ensure the preservation of order in the results.
-  • In some instances, the $\widehat{\text{Recall@5}}$ estimated by the URE scheme is inversely ordered compared to the actual Recall@5.
-This phenomenon can occur when the actual Recall@5 values are closely aligned, as is the case with the MF-based DR method at 0.353 and the MF-based AutoDebias method at 0.355.
-We contend that such a degree of deviation is within an acceptable range.
-  • When using $\widehat{\text{Recall@5}}$ to verify the debiasing performance, some debiasing methods perform worse than the Bias model, which encourages the development of further debiasing techniques.
-  • Based on the proposed URE, the performance of NegS is sub-optimal as expected in most instances, which implies that URE possesses potential merits in selecting truly effective debiasing approaches.
-In the LightGCN-based method on Yahoo!R3, the Recall@5 of NegS exceeds that of all other methods, which we attribute to the fact that these methods are not successfully debiasing in this case.
-  • There is an inconsistency in the ranking results based on $\widehat{\text{Recall@}K}$ metric with large and small $K$ values.
-This means we should develop different debiasing methods to accommodate Recall@$K$ with different $K$ on $D_{full}$.
-7 Related Work
-7.1 Debiasing Recommendation
-Recent recommendation debiasing strategies can be divided in two main categories: data-level debiasing and representation-level debiasing.
-At the data level, debiasing is always achieved by inverse propensity reweighting  (Schnabel et al.
-2016; Wang et al.
-2016; Li et al.
-2023c) and doubly robust learning (Dai et al.
-2022; Li et al.
-2023b; Li, Zheng, and Wu 2023).
-At the representation level, information bottleneck (Wang et al.
-2020), causal representation learning (Liu et al.
-2020; Yang et al.
-2023), and invariant representation learning (Wang et al.
-2022; Yang et al.
-2023; Pan et al.
-2023) are widely used.
-For evaluating the debiasing effectiveness of these methods, researchers often evaluate Recall@$\overline{K}$ performance (e.g., Recall@$\overline{5}$) of the learned recommendation models on publicly available randomly-exposed data (Chen et al.
-2023) — Coat, Yahoo!R3 and KuaiRand.
-However, we point out such an evaluation scheme on $D_{rand}$ may be problematic.
-7.2 Sampling Metrics
-The traditional evaluation scheme that calculates Recall@$\overline{K}$ metric on $D_{rand}$ can be thought of as a type of sampling metric.
-(Krichene and Rendle 2020) is the first work to highlight the inconsistency between sampled metrics on the sampled data and exact metrics on full data.
-It also proposes a method based on the sampled data to unbiasedly estimated Recall@$K$ metrics on full data.
-The simple estimation method suffers from the problem of large variance.
-Subsequent work (Li et al.
-2023a, 2020; Jin et al.
-2021; Dallmann, Zoller, and Hotho 2021) improves the estimation scheme and reduce the estimation variance.
-However, all the works mentioned above only focus on the case where the test set contains only one positive sample.
-In this way, the conclusions are not applicable to the debiasing task in that the test set contains multiple positive samples.
-Furthermore, there are no discussions about the effect of traditional evaluation scheme on the debiasing task.
-8 Conclusion And Future Work
-In this paper, we reviewed traditional evaluation scheme on the randomly-exposed datasets by comparing it with the gold standard evaluation on the fully-exposed datasets.
-According to the theoretical and empirical analyses, we revealed the inconsistency between them w.r.t.
-the values of $\overline{K}$ and $K$.
-We conducted an entrapment study, where a hard negative sampling trick without any debiasing techniques can achieve incredible debiasing performance by using the traditional evaluation scheme on the randomly-exposed datasets, further validating the traditional evaluation scheme on the randomly-exposed datasets may not reflect true debiasing performance and results in unreliable conclusions.
-Accordingly, we designed the URE scheme to unbiasedly estimate the true Recall performance of the fully-exposed datasets.
-Both theoretical and empirical analyses have validated the effectiveness of our proposal.
-In future work, we will investigate the unbiased estimation of other Top-$K$ metrics, like Precision and NDCG on the fully-exposed dataset to further extend our conclusions.
-Besides, improving URE from the perspective of bias-variance balancing is another interesting direction.
-We hope that the insight of this paper could bring new inspiration for debiasing recommendations.
-Appendix A Appendix
-A.1 Datasets
-We use two real-world datasets (Yahoo!R3 and KuaiRec) for our experiments.
-  • Yahoo!R3: This dataset is widely-used in debiasing literature which consists of a biased dataset and a randomly-exposed dataset collected by a random logging strategy.
-We follow (Chen et al.
-2021) and randomly sample 10% of the randomly-exposed data for debiasing training (required by AutoDebias and IPS), and take the remaining 90% as $D_{rand}$ for evaluating.
-The interactions between users and items are ratings ranging from one to five.
-The ratings larger than three are labeled as positive ($r_{u,i}$ = 1), otherwise negative ($r_{u,i}$ = 0).
-Notably, the Recall@$K$ metric is not available in Yahoo!R3, as it lacks a fully-exposed dataset.
-However, in our proposed URE scheme, we can use $D_{rand}$ to compute $\widehat{\text{Recall@}K}$ for unbiased estimation of Recall@$K$.
-  • KuaiRec: This dataset is the only existing fully-exposed dataset.
-We construct three types of datasets from this original fully-exposed data: biased data for training, fully-exposed data $D_{full}$ for evaluating, and randomly-exposed data $D_{rand}$ for evaluating.
-Concretely, we sample items for each user inversely proportional to their video duration and combine them with their ratings to obtain biased data.
-Interactions on the $k$-th shortest video are sampled with probability $p=e^{k/\lambda}/\sum_{k}e^{k/\lambda}$, $\lambda$ controls the degree of bias towards short videos.
-The remaining items and ratings is considered as the fully-exposed data ($D_{full}$) for evaluating.
-Following (Marlin and Zemel 2009), we construct randomly-exposed data ($D_{rand}$) by uniformly sampling $\overline{N}$ items for each user on $D_{full}$.
-The watch_ratio is binarized with threshold 2 following (Gao et al.
-2022a).
-That is, the video watching ratio larger than 2 is labeled as positive ($r_{u,i}$ = 1), otherwise negative ($r_{u,i}$ = 0).
-A.2 Proof of Theorem 3
-Proof.
-We first calculate the $E_{\mathcal{R}}[\text{Recall@}K]$.
-Since we consider all possible ranking, the possibility of the x positive samples contained in the top K is $\frac{C_{N^{+}}^{x}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$ where $C^{n}_{N}:=\frac{N!}{(N-n)!n!}$ is the number of combinations obtained by randomly sampling $n$ samples from $N$.
-Thus, we have:
-$E_{\mathcal{R}}\left[\text{Recall}@K\right]=\sum_{x=1}^{min(K,N^{+})}\frac{x}{N^{+}}\frac{C^{x}_{N^{+}}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$ (5)
-$=\sum_{x=1}^{min(K,N^{+})}\frac{xN^{+}K}{xN^{+}N}\frac{C^{x-1}_{N^{+}-1}C^{(K-1)-(x-1)}_{(N-1)-(N^{+}-1)}}{C_{N-1}^{K-1}}$ (6)
-$=\frac{K}{N}\left[\sum_{x=1}^{min(K,N^{+})}\left[C_{N^{+}-1}^{x-1}C_{(N-1)-(N^{+}-1)}^{(K-1)-(x-1)}/C^{K-1}_{N-1}\right]\right]$ (7)
-$=\frac{K}{N}\left[\ C_{N-1}^{K-1}/C_{N-1}^{K-1}\right]=\frac{K}{N}.$ (8)
-As for the $E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]$, we have:
-$E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]=E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right].$ (9)
-Since we consider all possible permutations, $E_{\mathcal{R},\mathcal{S}_{\overline{N}}}\left[\text{Recall@}\overline{K}|{N^{+}}^{\prime}\right]$ can be transformed into the scenario of randomly selecting $\overline{K}$ samples from $\overline{N}$ samples ($N^{+}$ positive samples among these $N$ samples) and determining how many of the selected $\overline{K}$ samples are positive.
-The problem is similar to Eq 5.
-Thus we have:
-$E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right]$ (10)
-$=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x}{{N^{+}}^{\prime}}\frac{C^{x}_{{N^{+}}^{\prime}}C^{\overline{K}-x}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}}^{\overline{K}}}\right]$ (11)
-$=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x\overline{K}{N^{+}}^{\prime}}{x{N^{+}}^{\prime}\overline{N}}\frac{C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{N-{N^{+}}^{\prime}}}{C_{N-1}^{\overline{K}-1}}\right]$ (12)
-$=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\frac{\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}-1}^{\overline{K}-1}}\right]$ (13)
-$=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\right]=\frac{\overline{K}}{\overline{N}}=\frac{K}{N}=E_{\mathcal{R}}[\text{Recall@}K].$ (14)
-A.3 Proof of Theorem 2
-Proof.
-It has been proven that when optimizing the BPR loss function using uniform negative sampling, the optimization objective of the model is AUC in the previous study (Rendle et al.
-2009).
-Therefore, we only need to prove that the BCE loss function commonly used in debiasing tasks is an upper bound of the BPR loss function.
-Specifically, the BCE loss function is:
-$\delta_{BCE}=\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\left[-\log(\sigma(f_{\theta}(u,i))-\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}\log(1-\sigma(f_{\theta}(u,j))\right],$ (15)
-where $\sigma(x)=1/(1+e^{-x})$, $f_{\theta}(u,i)$ denotes a model with trainable parameter $\theta$ for predicting the rating of item $i$ by user $u$, and $\mathcal{D}$ refers to the train dataset containing ($u$, $i$) pairs.
-BPR loss function:
-$\delta_{BPR}=\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log(\sigma(f_{\theta}(u,i)-f_{\theta}(u,j)),$ (16)
-where $\mathcal{I}$ denotes the item set.
-By transforming the BCE loss function, we can obtain:
-$\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\left[-\log(\sigma(f_{\theta}(u,i))-\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}\log(1-\sigma(f_{\theta}(u,j))\right]$ (17)
-$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log(\sigma(f_{\theta}(u,i)\cdot(1-\sigma(f_{\theta}(u,j)))))$ (18)
-$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\left[\frac{1}{1+e^{-f_{\theta}(u,i)}}\cdot\frac{e^{-f_{\theta}(u,j)}}{1+e^{-f_{\theta}(u,j)}}\right]$ (19)
-$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\frac{1}{1+e^{-f_{\theta}(u,i)}+e^{f_{\theta}(u,j)}+e^{-f_{\theta}(u,i)+f_{\theta}(u,j)}}$ (20)
-$\geq$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\frac{1}{1+e^{-(f_{\theta}(u,i)-f_{\theta}(u,j))}}=\delta_{BPR}.$ (21)
-Thus, we have proved that the commonly used BCE loss function is an upper bound of the BPR loss function.
-Therefore, optimizing the BCE loss function is equivalent to optimizing the upper bound of the BPR loss function, which equals optimizing AUC.
-Based on the aforementioned mentioned equivalence and referring to Theorem 1 and Theorem 3 from (Shi et al.
-2023), we derive that optimizing BCE loss enhanced with DNS($H$, $U$) approximates the optimization of the Recall@$K$ metric.
-Specifically, the larger the $H/U$ is, the Recall@$K$ metric with a larger $K$ on $D_{full}$ is optimized.
-∎
-A.4 Proof of Theorem 3
-Proof.
-Since Recall is only related to positive samples, we can only consider positive samples.
-Let’s assume that in a fully exposed scenario, there are a total of $N$ positive samples, out of which $M$ samples are ranked higher than the sample ranked at $K+1$ in Sorted ItemID in Figure 4.
-Assuming that we have randomly sampled $\overline{N}$ samples, which have $n$ positive samples and the rest are negatives.
-Since the sampling process is random, the probability of the number $m$ of positive samples that rank higher than the sample ranked at $K+1$ in Sorted ItemID can be calculated as $(C^{m}_{M}C^{n-m}_{N-M}/C^{n}_{N})$, thus we have:
-$E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall}@K}\right]$ $=E_{n^{\prime}}\left[E_{\mathcal{S}_{\overline{N}}}[m/n|n^{\prime}]\right]$ (22)
-$=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{m}{n^{\prime}}\frac{C^{m}_{M}C^{n^{\prime}-m}_{N-M}}{C^{n^{\prime}}_{N}}\right]$ (23)
-$=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{mn^{\prime}M}{mn^{\prime}N}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$ (24)
-$=E_{n^{\prime}}\left[\frac{M}{N}\sum_{{m=1}}^{min(M,n^{\prime})}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$ (25)
-$=M/N=\text{Recall}@K.$ (26)
-A.5 Compared Methods
-We utilize two representative recommendation models, Matrix Factorization (MF) (Koren, Bell, and Volinsky 2009) and graph neural network model LightGCN (He et al.
-2020), as the backbones.
-Based on the Theorem 2, We devise NegS, which optimizes the backbone model through BCE loss accompanied by the hard negative sampling, and compare NegS with following debiasing methods:
-  • IPS (Schnabel et al. 2016) weights the loss of each sample with an inversed propensity. We follow (Chen et al. 2021) and calculate the propensity with naive bayes.
-  • DR (Wang et al. 2019) integrates imputed errors and propensities while mitigating the impact of high propensity variance.
-  • AutoDebias (Chen et al. 2021) propose a meta-learning-based algorithm to learn debiasing parameters from a small set of randomly-exposed data.
-A.6 Metrics
-We use Recall metric to evaluate models for our following experiments.
-The Recall metric is the most important metric in the candidate generation stage of industrial recommendation systems.
-It is also a commonly used metric in research on debiasing recommendations.
-In future work, we will investigate the unbiased estimation of other Top-$K$ metrics, like Precision and NDCG on the fully-exposed dataset to further extend our conclusions.
-A.7 Setup
-All the models are optimized by Adam optimizer.
-We use grid search to find the best hyperparameters.
-The search space of learning rate and weight decay are [1e-4, 1e-3, 1e-2, 1e-1].
-For the proposed NegS, we tune $H/U$ and $U$ within the ranges of [0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1] and [50,100,200] respectively.
-For details on the dataset, see Appendix A.1.
-For details on the baseline, see Appendix A.5.
-For details on the matrics, see Appendix A.6.
+In contrast, the evaluation results of the traditional evaluation scheme exhibit a significant discrepancy from the true Recall@5, and it fails to ensure the preservation of order in the results.
+(2) When using $\widehat{\text{Recall@5}}$ to verify the debiasing performance, some debiasing methods perform worse than the orignal MF, which encourages the development of further debiasing techniques.
+Table 2. The statistics of datasets.
+Dataset ##\##Users ##\##Items Biased train test
+Yahoo!R3 15,400 1,000 311,704 48,600 ($D_{rand}$)
+KuaiRec 1411 3327 1,934,404 2,742,166 ($D_{full}$)
+6. Conclusion
+In this paper, we have unveiled the inconsistency between traditional evaluation scheme on the $D_{rand}$ and the gold standard evaluation on the $D_{full}$, which results in an inaccurate assessment of debiasing methods.
+To tackle this issue, we have introduced the URE scheme to unbiasedly estimate true Recall performance.
+Our proposal has been substantiated through both theoretical and empirical analyses, affirming its effectiveness.
+Acknowledgements.
+This work is supported by the National Key Research and Development Program of China (2022YFB3104701) and the National Natural Science Foundation of China (62272437).